In the gravimetric analysis of iron, hydroxide may be added to a solution containing Fe2+ to

precipitate a gelatinous mess which is ignited to form Fe2O3. If 0.2864 grams of Fe2O3 were formed

from the ignition process, how many grams of FeCO3 were contained in the original sample that

was analyzed?
1
Expert's answer
2020-09-29T06:53:42-0400

1) FeCO3+2NaOH=Fe(OH)2+Na2CO3

2) 4Fe(OH)2+O2=2Fe2O3+4H2O


M(Fe2O3)=159.6900 g/mol

n (Fe2O3)=m/M=0.2864/159.6900=0.0018 mol


In the second reaction: n (Fe(OH)2)=2 x n (Fe2O3)=0.0018 x 2 = 0.0036 mol


In the first reaction: n (FeCO3) = n (Fe(OH)2) = 0.0036 mol


M (FeCO3) = 115.8540 g/mol

m (FeCO3) = M x n = 115.8540 x 0.0036 = 0.4156 g